# Generative AI and Firm Productivity:

### v2

@@ 4.3 Aggregate Productivity Gains Across Workflows
In this section, we aggregate productivity gains across workflows to estimate the AI-driven incremental value per consumer observed in our experiments. Specifically, we focus on the four workflows with positive treatment effects in sales, excluding the Google Advertising Title experiment, which serves as a testable bad case and can be readily improved in future iterations.
@@
1 Pre-sale Service Chatbot 0.274 (upper-bound)
@@
2 Search Query Refinement 0.0648 40.6 2.63
3 Product Description 0.0104 52.1 0.54
4 Marketing Push Message 0.000402 365.0 0.15
@@
Table 6 summarizes the key variables used in our calculations. Column 1 reports the AI-driven incremental value per consumer for each experimental workflow, reflecting the estimated average treatment effects (the absolute sales lift per consumer) reported in Column 1 of Table 4. Because each experiment had a different duration, Column 2 presents the time multiplier used to extrapolate these effects to an annual horizon. [^18] We assume that each experiment’s duration reflects the typical interval between treatment opportunities for a representative consumer on the platform. For example, a representative consumer is assumed to interact with the Pre-sale Service Chatbot about once every two months, whereas they could receive a Marketing Push Message daily. For instance, the Pre-sale Service Chatbot experiment spans two months, yielding a time multiplier of six (i.e., $12/2$). For each workflow, we obtain the annualized AI-driven incremental value per consumer in Column 3 by multiplying the estimated effect in Column 1 by the corresponding time multiplier in Column 2. This calculation assumes that the treatment effects observed during the experimental period remain constant over time, abstracting from potential amplification (e.g., greater engagement and purchases as GenAI-generated content enhances consumer satisfaction and platform loyalty) or attenuation (e.g., consumer dissatisfaction and product returns due to potential mismatches between AI-generated content and actual product characteristics).
@@
Finally, we aggregate the annualized estimates across the four workflows to obtain the total annual incremental value per consumer attributable to GenAI (see “Total” in Column 3). This aggregation assumes that effects across workflows are linearly additive and abstracts from potential cross-workflow interactions, such as cannibalization across touchpoints or expansion through synergies. Based on the four GenAI applications with positive sales effects, we estimate an annual incremental value of up to \$4.96 per consumer, which decreases to \$4.63 when using the lower-bound estimate from the Pre-sale Service Chatbot experiment. These effects represent roughly 5.5-6% of the per-user revenue growth observed in global e-commerce between 2023 and 2024 [59], highlighting the economic significance of these gains relative to broader industry trends.
@@
It is worth noting that these estimates capture only a partial and time-dependent view of the firm’s efforts to scale up GenAI across workflows. While in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it was deployed in more than 40 applications and by 2025 in over 60. This rapid expansion is also reflected in the growth of API calls to the platform’s GenAI tools: in mid-2024, AI-related API requests averaged over 50 million per day, rising to more than 1 billion per day by mid-2025—a twentyfold increase. Hence, our point estimates should be interpreted with caution and in light of these rapid adoption trends, which nonetheless indicate that the firm anticipates substantial value from GenAI.
@@
Taken together, these results reveal sizable gains in targeted workflows and measurable contributions to overall platform sales, with substantial potential as GenAI applications diffuse across use cases and models are further refined for domain-specific tasks. This pattern aligns with [1], who emphasize that the aggregate productivity effects of new general-purpose technologies materialize gradually as complementary investments and organizational adaptations accumulate.
@@ 5 Heterogeneous Treatment Effects
To further shed light on the mechanism behind our results, we examine heterogeneity in treatment effects across sellers, buyers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, one would expect relatively larger gains among participants with lower baseline capabilities—namely, smaller and newer sellers, less experienced buyers with limited platform engagement, and products in the long tail of sales or in less concentrated categories. We classify each dimension into “high” and “low” groups based on pre-experiment characteristics that capture scale, experience, or market position. For sellers, the classification is based on size and tenure on the platform. For buyers, we use measures of online shopping experience and purchasing intensity. For products, we rely on indicators of category concentration, sales volume, and relative price. These groupings allow us to assess whether GenAI adoption disproportionately benefits smaller or newer sellers, less sophisticated buyers, and products in the long tail or in less concentrated categories, thereby revealing how the technology shapes outcomes across different segments of the marketplace. Note that, depending on the context and data availability, some heterogeneity analyses cannot be implemented for certain workflows. We find consistent results across seller and buyer characteristics: disadvantaged groups, such as small sellers and inexperienced buyers, experience greater gains from GenAI-powered workflows. In contrast, the effects across product groups are more context-dependent.
@@ 5.1 Heterogeneous Effects Across Sellers
The platform hosts a highly diverse population of sellers, varying substantially in firm characteristics. We focus on sellers’ size and sophistication to investigate whether the productivity potential of GenAI vary between big versus small sellers. We classify sellers into high and low groups based on three pre-experiment metrics: annual sales value (Annual Sales), years of operation on the platform (Operation Years), and the number of sub-accounts linked to the seller’s online store (# of Sub-Accounts). For each metric, sellers in the low group are generally small sellers, defined as meeting the following pre-experiment criteria: (i) accounting for the bottom 50% cumulative share of total sales when sellers are ranked by annual sales; (ii) having operated on the platform for fewer than five years; [^19] The platform closely monitors seller tenure, with five years of continuous operation serving as an important performance indicator. or (iii) maintaining fewer than three sub-accounts for their online store. [^20] According to the advice of the platform’s internal staff, small online stores are typically run by individuals or operate as mom-and-pop businesses, often with no more than two sub-accounts. Stores that have more sub-accounts generally employ additional staff to assist with tasks related to store operations, suggesting that these are more likely to be larger sellers. Table 7 presents the summarized results on seller heterogeneity, reporting only the percent change and significance levels for brevity. Additional results and details can be found in Appendix D.
Table 7: Summary of Heterogeneous Treatment Effects Across Sellers
(1) (2) (3) (4) (5) (6)
Business Dependent Annual Sales Operation Years # of Sub-Accounts
Workflow Variable High Low High Low High Low
1 Search Query Sales 2.18% 3.68%** 2.28% 3.19%** 0.97% 3.48%**
Refinement Conversion 0.21% 1.69%*** 1.24% 1.01%* 0.88% 1.20%**
2 Marketing Sales 1.8% 1.4% 5.1% 0.6% 6.8% -1.0%
Push Message Conversion 3.3% 2.7% 0.9% 3.8%** -0.1% 4.8%**
3 Google Sales -5.1% -4.4% -5.4% -4.4% -5.9% -4.5%
Ad Titles Conversion -6.5% -0.3% -2.3% -3.8% -4.7% -3.2%
  • 1
    We classify sellers into high and low groups based on three different pre-experiment metrics. Low-group sellers are generally small sellers, as defined by meeting the following pre-experiment criteria: (1) accounting for the bottom 50% cumulative share of total sales when sellers are ranked by annual sales; (2) having operated on the platform for fewer than five years; or (3) maintaining fewer than three sub-accounts for their online store.
  • 2
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
  • 3
    We report the % change. % change is calculated by dividing the treatment effect by the control group average. The asterisks notation indicates the significance of treatment effect. *** p<<0.01, ** p<<0.05, * p<<0.1.
  • 4
    The heterogeneity analysis of different seller types was not applicable for experiments of Pre-sale Service Chatbot and Product Description as these two were conducted on platform self-sold products, which involved only a few platform-operated sellers, thereby lacking variation in seller characteristics for meaningful examination of seller heterogeneity.
Search Query Refinement.
Table 7 shows that small sellers experience significant gains in both sales and conversion rates (Columns 2, 4, and 6), whereas effects for large sellers tend to be smaller and statistically insignificant (Columns 1, 3, and 5). GenAI enhances the search algorithm’s ability to translate queries by understanding consumer intent and refining queries to improve semantic accuracy and clarity. This refinement allows the search engine to retrieve products more relevant to consumer needs, thereby reducing search frictions, improving matching efficiency, and ultimately increasing purchases for small sellers. These findings are consistent with prior work showing that lowering search costs or improving search quality fosters long-tail dynamics and expands market share for smaller and niche sellers [14, 5, 65, 69].
Marketing Push Message.
A similar pattern, but only with respect to conversion rates, is observed in the platform’s advertising function, where only small sellers show significant gains. A likely explanation is that human drafters, constrained by time and attention, tend to reference or replicate the product characteristics of top sellers when composing marketing content. By contrast, GenAI enables the large-scale generation of customized marketing messages across a broad product portfolio, enhancing content differentiation for small sellers’ offerings and stimulating consumer purchases. These findings align with [60], who show that personalized recommendation systems disproportionately benefit smaller and less established sellers.
Google Advertising Title.
In this experiment, heterogeneous treatment effects across seller types are statistically insignificant for both high and low groups.
Other Workflows.
We cannot assess seller heterogeneity in the Pre-sale Service Chatbot and Product Description experiments, as both were conducted exclusively on platform self-sold products with only a handful of platform-operated sellers, leaving insufficient variation for meaningful analysis. In the case of Product Description, as well as for experiments where granular data are unavailable (e.g., Chargeback Defense), the platform’s internal data science team provides descriptive evidence from third-party sellers to offer anecdotal insights into potential heterogeneity in sellers’ responses to GenAI. Specifically, third-party sellers are divided into quartiles based on pre-experimental annual sales. The share of products with no or limited descriptions is 1.2 times higher among bottom-quartile (small) sellers relative to top-quartile (large) sellers, while the share of unaddressed chargeback disputes is 1.7 times higher. These figures suggest that small sellers are particularly constrained by shortages of talent and resources, lacking the capacity to perform or outsource functions such as customer service, product content generation, or compliance management. Taken together, these patterns provide anecdotal support for the view that smaller sellers stand to gain disproportionately more from GenAI.
Overall, our findings indicate that the productivity gains from GenAI are disproportionately larger for small sellers, characterized by lower transaction value, shorter operational histories, and fewer sub-accounts. This pattern is consistent with prior research on GenAI adoption by individual workers, which shows that low-skilled and less experienced labor—often disadvantaged by earlier technological advances—derive greater benefits from GenAI [15, 43, 21, 54, 26, 22, 56]. Our results extend this insight to the domain of entrepreneurship on online retail platforms.
5.2 Heterogeneous Effects Across Consumers
Beyond seller heterogeneity, little research has examined the impact of GenAI on different types of consumers. Our analysis focuses on the differential effects of the technology on experienced versus inexperienced consumers. We classify consumers into high and low groups based on three pre-experiment indicators of online shopping experience: years since registration on the platform (Registered Years), number of login days during the 30 days prior to the experiment (Past Login Days), and total expenditure during the same period (Past Sales). For each indicator, consumers in the low group are defined as relatively inexperienced if they meet the following criteria: (i) a registration duration below the platform-wide median; (ii) login days below the platform-wide median; or (iii) total purchases below the platform-wide median. Our results show significant heterogeneity in GenAI’s effects across these consumer segments. Summary results are presented in Table 8, with additional details reported in Appendix E.
Table 8: Summary of Heterogeneous Treatment Effects Across Consumers
(1) (2) (3) (4) (5) (6)
Business Dependent Registered Years Past Login Days Past Sales
Workflow Variable High Low High Low High Low
1 Pre-sale Sales 13.7%** 22.4%** 15.0%** 18.5%** 8.6% 25.9%***
Serv Chatbot Conversion 19.8%*** 26.1%*** 17.1%*** 29.6%*** 17.6%*** 25.4%***
2 Search Query Sales 1.27% 5.01%** 0.63% 8.16%*** -0.56% 7.46%***
Refinement Conversion -0.09% 2.79%*** 0.43% 2.32%*** 0.44% 1.93%***
3 Product Sales 1.09% 3.06%*** -1.02% 6.24%*** 0.03% 5.63%***
Description Conversion 1.10%* 1.39%** 0.65% 2.06%*** 1.05%* 1.59%**
4 Marketing Sales -0.9% 6.6% 2.5% 0.8% -5.9% 20.6%***
Push Message Conversion 1.8% 5.0%** 0.7% 4.8%*** -2.2% 14.2%***
  • 1
    We classify consumers into high and low groups and capture their online shopping experience using three pre-experiment indicators. Consumers in the low group are considered relatively inexperienced, as determined if they meet the following criteria: (1) a registration duration below the platform-wide median; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
  • 2
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
  • 3
    We report the % change. % change is calculated by dividing the treatment effect by the control group average. The asterisks notation indicates the significance of treatment effect. *** p<<0.01, ** p<<0.05, * p<<0.1.
  • 4
    The heterogeneity analysis of different consumer types is not applicable for the workflow of Google Advertising Title because the platform cannot access detailed consumer demographic data from Google.
Pre-sale Service Chatbot.
Table 8 shows that although both consumer groups benefit from the intervention, the gains are relatively larger for inexperienced consumers—namely, those with shorter registration histories, fewer prior login days, and lower past spending (Column 2, 4 and 6). One explanation is that experienced online shoppers are generally able to locate and evaluate product information independently, even when customer support is limited to pre-programmed auto responses (as in the control condition). By contrast, less experienced users rely more heavily on customer service for product and seller information to inform their decisions. As a result, GenAI-based customer support enables these consumers to navigate the platform more effectively and make better purchase choices.
Search Query Refinement.
In this experiment, increases in sales and conversion rates were significant and pronounced only for less experienced consumers, while the effects for more experienced consumers were not statistically significant. This pattern likely reflects the challenges faced by inexperienced consumers, who generally possess weaker information-search skills and less familiarity with the online search environment. Consequently, they may struggle to articulate their needs effectively through query-based searches and thus derive greater benefits from GenAI-enhanced search.
Product Description.
Augmenting product descriptions with GenAI-generated content led to substantial increases in consumer purchases, particularly among inexperienced consumers, as reflected in the significant sales gains for this group. A plausible explanation is that inexperienced consumers often lack domain knowledge and access to alternative information sources relative to experienced users, and therefore rely more heavily on detailed descriptions to guide their decisions. These findings suggest that GenAI can help bridge informational gaps and improve the shopping experience for less experienced consumers.
Marketing Push Message.
In this experiment, the results remain consistent with the previous findings. When GenAI is used to enhance the personalization and differentiation of marketing messages across a broad product portfolio, the perceived relevance of the content increases, making it more likely to capture the attention of less experienced consumers and support their purchase decisions.
Google Advertising Title.
This experiment is not suitable for consumer heterogeneity analysis, due to the fact that the it was conducted off-site at the product level, preventing the platform from accessing detailed consumer demographic data from Google.
To sum up, we find that inexperienced consumers—those with shorter registration histories, lower login frequencies, and lower spending levels—benefit more from GenAI than their more experienced counterparts. This result is consistent with prior evidence that GenAI can support vulnerable groups such as low-skilled workers [15, 54, 56], and it extends this insight to the consumer domain by showing that GenAI disproportionately benefits less sophisticated consumers more. It also aligns with existing research indicating that enhanced e-commerce technologies—such as improved information provision, search refinement, and personalized recommendation systems—deliver greater benefits to consumers who are older, newer to the platform, or have lower purchasing power [33, 60], while extending these patterns to the new context of GenAI.

### v3

@@ 4.3 Aggregate Productivity Gains Across Workflows
In this section, we aggregate productivity gains across workflows to estimate the AI-driven incremental value per consumer observed in our experiments, explicitly accounting for differences in the frequency of consumer interactions across workflows. Specifically, we focus on the four workflows with positive treatment effects in sales, excluding the Google Advertising Title experiment, which serves as a testable bad case and can be readily improved in future iterations.
@@
Pre-sale Service Chatbot 0.274 (upper-bound)
@@
Search Query Refinement 0.0648 40.6 2.63
Product Description 0.0104 52.1 0.54
Marketing Push Message 0.000402 365.0 0.15
@@
Table 6 summarizes the key variables used in our calculations. Column 1 reports the AI-driven incremental value per consumer for each experimental workflow, reflecting the estimated average treatment effects (the absolute sales lift per consumer) reported in Column 1 of Table 4. Because each experiment had a different duration, Column 2 presents the time multiplier used to extrapolate these effects to an annual horizon. [^19] We assume that each experiment’s duration reflects the typical interval between treatment opportunities for a representative consumer on the platform. For example, a representative consumer is assumed to interact with the Pre-sale Service Chatbot about once every two months, whereas they could receive a Marketing Push Message daily. For instance, the Pre-sale Service Chatbot experiment spans two months, yielding a time multiplier of six (i.e., $12/2$). For each workflow, we obtain the annualized AI-driven incremental value per consumer in Column 3 by multiplying the estimated effect in Column 1 by the corresponding time multiplier in Column 2. This calculation assumes that the treatment effects observed during the experimental period remain constant over time, abstracting from potential amplification (e.g., greater engagement and purchases as AI-generated content enhances consumer satisfaction and platform loyalty) or attenuation (e.g., consumer dissatisfaction and product returns due to potential mismatches between AI-generated content and actual product characteristics).
@@
Finally, we aggregate the annualized estimates across the four workflows to obtain the total annual incremental value per consumer attributable to GenAI (see “Total” in Column 3). This aggregation assumes that effects across workflows are linearly additive and abstracts from potential cross-workflow interactions, such as cannibalization across touchpoints or expansion through synergies. Based on the four GenAI applications with positive sales effects, we estimate an annual incremental value of up to \$4.96 per consumer, which decreases to \$4.63 when using the lower-bound estimate from the Pre-sale Service Chatbot experiment (human agents vs. GenAI assistance equipped with human escalation). These effects represent roughly 5.5-6% of the per-user revenue growth observed in global e-commerce between 2023 and 2024 [undefaaf], highlighting the economic significance of these gains relative to broader industry trends.
@@
It is worth noting that these estimates capture only a partial and time-dependent view of the firm’s efforts to scale up GenAI across workflows. While in 2023 the platform applied GenAI to only a handful of workflows, by 2024 it was deployed in more than 40 applications and by 2025 in over 60. This rapid expansion is also reflected in the growth of API calls to the platform’s GenAI tools: in mid-2024, AI-related API requests averaged over 50 million per day, rising to more than 1 billion per day by mid-2025, representing a twentyfold increase. Hence, our point estimates should be interpreted with caution and in light of these rapid adoption trends, which nonetheless indicate that the firm anticipates substantial value from GenAI.
@@
Taken together, these results reveal sizable gains in targeted workflows and measurable contributions to overall platform sales, with substantial potential as GenAI applications diffuse across use cases and models are further refined for domain-specific tasks. This pattern aligns with [undef], who emphasize that the aggregate productivity effects of new general-purpose technologies materialize gradually as complementary investments and organizational adaptations accumulate.
@@ 5 Heterogeneous Treatment Effects
A central question in online retail is which platform participants benefit most from GenAI-driven improvements. To address this, we examine heterogeneity in treatment effects across buyers, sellers, and products. If GenAI primarily reduces frictions on both the demand and supply sides, one would expect relatively larger gains among participants with lower baseline capabilities—namely, less experienced buyers with limited platform engagement, smaller and newer sellers, and products in the long tail of the sales distribution or in less concentrated categories.
@@
For each dimension, we classify observations into “high” and “low” groups based on pre-experiment characteristics that capture experience, scale, or market position. For buyers, we use measures of prior online shopping experience and purchasing intensity. For sellers, we classify firms based on size and tenure on the platform. For products, we rely on indicators of category concentration, sales volume, and relative price. These groupings allow us to assess whether GenAI adoption disproportionately benefits less sophisticated buyers, smaller or newer sellers, and products in the long tail, thereby shedding light on how GenAI reshapes outcomes across segments of the marketplace.
We find consistent evidence of heterogeneity on the demand side: buyers with lower prior spending and less platform experience benefit more from GenAI-powered workflows. On the supply side, the estimated effects point to larger gains for smaller and newer sellers, although these estimates are less precise. Finally, heterogeneity across product groups is more context-dependent. We note that, due to data limitations and workflow-specific features, some heterogeneity analyses cannot be implemented for all applications.
5.1 Heterogeneous Effects Across Consumers
We classify consumers into high- and low-experience groups based on three pre-experiment indicators of online shopping experience: total expenditure over the period preceding the experiment (Past Purchases), number of login days in the 30 days prior to the experiment (Past Login Days), and years since registration on the platform (Registered Years). For each indicator, consumers in the low group are defined as relatively inexperienced if they fall below the median of the corresponding distribution. We estimate treatment effects separately for high- and low-experience consumers and test for differences in percentage treatment effects across groups using Wald tests implemented via seemingly unrelated estimation.
Figure 3: Heterogeneous Treatment Effects on Sales Across Consumers
Chatbot Refer to caption
Search Refer to caption
Description Refer to caption
Push Refer to caption
Notes: The bars indicate the % change in treatment effects for the high (blue) and low (red) groups. The p-values are from seemingly unrelated estimations testing the equality of % treatment effects across the two groups.
Results on sales are summarized in Figure 3, with regression results and additional evidence on conversion rates and cart values reported in Appendix D. [^20] The Google Advertising Title experiment is not suitable for consumer-level heterogeneity analysis, as it was conducted off-platform at the product level, preventing access to detailed consumer characteristics. Overall, we find statistically significant heterogeneity in GenAI’s effects across consumer segments. We begin by analyzing heterogeneity by past purchases, our preferred proxy for consumers’ prior experience on the platform. Across workflows, GenAI-driven gains are significant and systematically larger for consumers with fewer past purchases, while the effects for consumers with more purchase experience are smaller and insignificant. Further analysis confirms that the differences between these two consumer groups are statistically significant. We find qualitatively similar patterns when stratifying consumers by past login activity and platform tenure, although the differences along these dimensions are generally less pronounced.
Our results indicate that the technology disproportionately benefits less experienced and less sophisticated users, aligning with prior evidence that GenAI can disproportionately benefit vulnerable groups, such as low-skilled workers [undefn, undefaaa, undefaac], and extend this insight to the consumer domain. The results are also consistent with earlier work demonstrating that improvements in e-commerce technologies, such as enhanced information provision, search refinement, and personalization, generate greater benefits for consumers who are newer to the platform or have lower purchasing power [undefaag, undefaf].
A plausible explanation for the observed buyer heterogeneity is that less experienced consumers face greater informational and search frictions throughout the purchase journey. With limited domain knowledge and fewer alternative information sources, they rely more heavily on customer service and product descriptions to acquire relevant information. They also exhibit weaker search skills and less familiarity with online search environments, making GenAI-enhanced search particularly valuable for articulating their needs and identifying relevant products. In addition, GenAI-driven message personalization increases the perceived relevance of marketing content, which is especially effective in supporting the purchase decisions of less experienced consumers.
@@ 5.2 Heterogeneous Effects Across Sellers
We examine heterogeneity in GenAI’s effects across sellers by focusing on differences in seller size and sophistication. Sellers are classified into high and low groups based on three pre-experiment characteristics: annual sales volume (Annual Sales), years of operation on the platform (Operation Years), and the number of sub-accounts associated with the seller’s store (# of Sub-Accounts). Sellers in the low group correspond to smaller and less established firms, defined by lower sales, shorter platform tenure, or fewer sub-accounts. [^21] Sellers are classified into the low group if they (i) account for the bottom 50% of cumulative annual sales, (ii) have operated on the platform for fewer than five years (the platform treats five years as a key performance benchmark), or (iii) maintain fewer than three sub-accounts (following platform guidance, small stores are typically individually operated or mom-and-pop businesses, whereas stores with more sub-accounts generally employ additional staff). Platform self-operated sellers are classified as large. We then decompose consumer purchases into transactions involving high- and low-group sellers and estimate treatment effects separately for each group. Differences in percentage treatment effects are formally tested using Wald tests implemented via seemingly unrelated estimation.
Summary results are reported in Figure 4, with regression results and additional details on conversion rates and cart values provided in Appendix E. Overall, we find limited statistical evidence of differential effects across seller types. While point estimates often suggest larger gains for smaller and less established sellers, most differences between high- and low-group sellers are statistically insignificant, reflecting imprecise estimates and limited power. This pattern holds across the workflows for which seller-level heterogeneity can be assessed. [^22] We cannot assess seller heterogeneity in the Pre-sale Service Chatbot and Product Description experiments, as both were conducted exclusively on platform self-sold products with only a handful of platform-operated sellers, leaving insufficient variation for meaningful analysis.
Figure 4: Heterogeneous Treatment Effects on Sales Across Sellers
Search Refer to caption
Push Refer to caption
Google Refer to caption
Notes: The bars indicate the % change in treatment effects for the high (blue) and low (red) groups. The p-values are from seemingly unrelated estimations testing the equality of % treatment effects across the two groups.